# A Spatial Gap in the Sky Distribution of Fast Radio Burst Detections Coinciding with Galactic Plasma Overdensities

### v1

@@ 6 Conclusion
Our analysis shows that FRB detections in CHIME/FRB Catalog 2 exhibit a two-dimensional dependence on Galactic coordinates. The most striking feature is the complete absence of CHIME/FRB detections in a patch of the sky that coincides with the Cygnus X region. The analyses discussed in §35 demonstrate that this detection gap, along with other regions of reduced FRB detections within the Galactic plane, correspond with known H II regions. Based on the discussion presented in §5, we favor the interpretation that turbulence along these lines of sight induces large degrees of scatter-broadening that totally smear FRBs out of detectability within the CHIME band. We note that measurements of $\tau_{\rm sc}$ in CHIME/FRB Catalog 2 and derived $\tau_{\rm sc}$ from non-detections could be used to gauge the accuracy of existing electron-density models of the Milky Way Galaxy.
@@
Future CHIME/FRB measurements of $\tau_{\rm sc}$ will increase the surface density of FRB detections at low frequencies. Such an increase will likely yield additional detection gaps associated with Galactic H II regions that have smaller angular extents. These same data sets will also allow for data-driven constraints on the Galactic electron density models themselves , through all-sky estimates of the maximum $\tau_{\rm sc}$ as estimated for the Cygnus X region in §4.4. Although the nature of FRB classification makes similar constraints on DM more complicated, phenomenological models can be used to gauge the likely maximum values for the contribution of the Milky Way to DM (e.g., A. M. Cook et al., 2023). We therefore expect that future CHIME/FRB catalogs will provide the means to strengthen and expand upon the inferences made in our work.
@@ ACKNOWLEDGEMENTS
We acknowledge that CHIME is located on the traditional, ancestral, and unceded territory of the Syilx/Okanagan people. We are grateful to the staff of the Dominion Radio Astrophysical Observatory, which is operated by the National Research Council of Canada. CHIME operations are funded by a grant from the NSERC Alliance Program and by support from McGill University, University of British Columbia, and University of Toronto. CHIME was funded by a grant from the Canada Foundation for Innovation (CFI) 2012 Leading Edge Fund (Project 31170) and by contributions from the provinces of British Columbia, Québec and Ontario. The CHIME/FRB Project was funded by a grant from the CFI 2015 Innovation Fund (Project 33213) and by contributions from the provinces of British Columbia and Québec, and by the Dunlap Institute for Astronomy and Astrophysics at the University of Toronto. Additional support was provided by the Canadian Institute for Advanced Research (CIFAR), the Trottier Space Institute at McGill University, and the University of British Columbia (UBC). This research has made use of the VizieR catalogue access tool, CDS, Strasbourg, France (DOI : 10.26093/cds/vizier). The original description of the VizieR service was published in 2000, A&AS 143, 23. We thank Dr. Duncan Lorimer, Dr. Liam Connor and Dr. Namir Kassim for thoughtful questions and discussions that informed the analysis presented in this paper.
@@ Appendix A Literature Scattering Values near Cygnus X
Table 1: Table with literature values of scattering and angular broadening towards Cygnus X, from background pulsars, FRBs, quasars, and Cygnus X-3. References: [1] W. C. Jing et al. (2025), [2] The CHAMPSS Collaboration et al. (2025), [3] D. J. Zhou et al. (2022), [4] A. L. Fey et al. (1989), [5] A. L. Fey et al. (1991), [6] L. A. Molnar et al. (1995)
Source $\ell$ (deg) $b$ (deg) $\theta$ (mas) $\tau_{\mathrm{1\,GHz}}$ (ms) $\tau_{\mathrm{600\,MHz}}$ (ms) Reference
PSR J2021+4024g 78.18 2.11 8.21e+01 6.33e+02 [1]
PSR J2030+3818g 77.45 -0.51 7.80 6.02e+01 [1]
PSR J2030+3944g 78.60 0.30 1.14e+02 8.80e+02 [1]
PSR J2046+4253g 83.02 -0.26 1.27e+02 9.80e+02 [1]
PSR J2052+4421g 84.84 -0.17 9.50e+01 7.33e+02 [1]
PSR J2108+5001 91.20 1.47 1.91 14.7 [2]
FRB 20210705 76.03 2.71 1.19e+01 9.20e+01 [3]
@@ B.1 Catalog
Table 2: A full list of FRB sightlines intersecting H II regions from the WISE (L. D. Anderson et al., 2014) and HH14 (L. G. Hou & J. L. Han, 2014b) catalogs.
TNS Name (l, b) $\tau_{\rm sc}$ DM Name (l, b) D $\theta_{\rm IR}$ Catalog
(deg) (ms at 1 GHz) (pc cm-3) (deg) (kpc) (deg)
FRB20200306C 15.6, 28 337.2 6.4, 22.9 5.19 WISE
FRB20230511B 15.7, 23.1 191.6 6.4, 22.9 5.19 WISE
Table 2: Continued.
TNS Name (l, b) $\tau_{\rm sc}$ DM Name (l, b) D $\theta_{\rm IR}$ Catalog
(deg) (ms at 1 GHz) (pc cm-3) (deg) (kpc) (deg)
FRB20181119B 67.8, 1 2.307 ± 2.3 608.5 68.2, 1.1 0.22 WISE
FRB20210817B 90.1, 3.6 483.1 90.2, 2.3 HH14
FRB20221109A 93.6, -6.4 1580 94.4, -5.5 HH14
FRB20201225A 94.3, 18.6 145.6 95.4, 16.8 HH14
FRB20190211B 94.4, 16.1 260.7 95.4, 16.8 HH14
FRB20230727C 94.9, 17 700.8 95.4, 16.8 HH14
FRB20210503A 95.1, 14.8 388.3 95.4, 16.8 HH14
FRB20200514D 95.4, 18.7 0.252 ± 0.25 1819 95.4, 16.8 HH14
FRB20211117F 95.7, 14.7 1021 95.4, 16.8 HH14
FRB20221206I 96.8, 18.7 707.1 95.4, 16.8 HH14
FRB20210105A 96.8, 7.8 2470 98.5, 8 HH14
FRB20210129A 97.7, 16.1 632.2 95.4, 16.8 HH14
FRB20180918A 98.1, 17.1 1454 95.4, 16.8 HH14
FRB20230426F 99.4, 1.4 0.286 ± 0.29 1360 S131; IC1396 99.5, 3.8 0.91 ± 0.57 1.26 WISE
FRB20210810B 101.4, 4.4 1.101 ± 1.1 639.6 S131; IC1396 99.5, 3.8 0.91 ± 0.57 1.26 WISE
FRB20220822A 102, 8.6 6.995 ± 7 1174 103.1, 9.6 HH14
FRB20221214C 102.3, 8.4 8.46 ± 8.5 409.5 103.1, 9.6 HH14
FRB20210507B 104.3, 10.9 415.5 103.1, 9.6 HH14
FRB20211103D 104.4, 7.5 0.689 ± 0.69 711.1 105.6, 7.8 HH14
FRB20191213C 105.7, 9.2 0.982 ± 0.98 571.3 105.4, 9.9 HH14
FRB20200711E 106.5, 10.1 691.6 105.4, 9.9 HH14
FRB20210103A 106.7, -0.2 1.881 ± 1.9 746.2 S142 107, -0.8 3.52 ± 0.47 0.37 WISE
FRB20230102D 107.2, 6 983 107.9, 5.8 HH14
S141 107.9, 5.6 0.55 WISE
S150 108.5, 6.4 1.34 ± 0.47 0.82 WISE
106.8, 5.3 HH14
FRB20200917B 108.3, 2.5 0.409 ± 0.41 1198 109.9, 2.1 HH14
FRB20230409B 112.3, 4.6 0.587 ± 0.59 1459 111.9, 4.1 HH14
FRB20210603D 114.7, 0.8 502.2 114.6, 0.2 HH14
FRB20230316H 117.3, 3.9 496.1 S171 118.2, 5.4 1.39 WISE
FRB20220411A 117.5, 8 0.323 ± 0.32 507.4 S171 118.2, 5.4 1.39 WISE
FRB20210309F 118.5, 6.5 547 S171 118.2, 5.4 1.39 WISE
FRB20200828D 118.9, 7.7 0.96 ± 0.96 1517 S171 118.2, 5.4 1.39 WISE
FRB20190912D 119.3, 3.9 786.2 S171 118.3, 4.9 0.84 ± 0.31 0.73 WISE
S171 118.2, 5.4 1.39 WISE
FRB20220218A 120.1, 5.4 594.5 S171 118.2, 5.4 1.39 WISE
FRB20200806B 121.8, 0.8 2766 121.3, 0.7 HH14
FRB20220330A 124.9, 0.9 0.367 ± 0.37 664.9 125, 0.9 0.12 WISE
FRB20220814A 127.5, 1 0.919 ± 0.92 1431 127, 0.8 HH14
127.6, 1.1 0.2 WISE
FRB20190302A 133.2, 0.2 11.28 ± 11 1034 S190; W3 133.5, 0.8 0.46 WISE
FRB20200416A 134.8, 1.2 0.702 ± 0.7 538.6 134.8, 0.9 HH14
W4; S190 135.8, 0.9 0.69 WISE
W4; S190 135, 0.7 3.31 ± 0.35 0.87 WISE
FRB20181014A 137.2, 4.2 0.478 ± 0.48 1314 138.1, 4.1 HH14
FRB20191231B 138, 3.9 703.2 138.1, 4.1 HH14
Table 2: Continued.
TNS Name (l, b) $\tau_{\rm sc}$ DM Name (l, b) D $\theta_{\rm IR}$ Catalog
(deg) (ms at 1 GHz) (pc cm-3) (deg) (kpc) (deg)
FRB20200708A 138.2, -1.1 7.998 ± 8 842.5 139.7, -1.2 1.52 WISE
FRB20230218B 138.3, 4.1 590.9 138.1, 4.1 HH14
FRB20191209B 138.9, 2.9 910.7 139.6, 2.5 3.81 ± 0.34 0.38 WISE
FRB20221129E 140.8, 2.7 1.709 ± 1.7 2293 141.7, 2.8 HH14
140.8, 3.1 HH14
140.8, 3.1 0.21 WISE
140.6, 1.9 HH14
FRB20210512A 141.4, 4.5 0.345 ± 0.34 778.7 140.8, 3.1 HH14
FRB20190325C 142, 3.5 797.9 141.7, 2.8 HH14
140.8, 3.1 HH14
FRB20210207E 142.6, 2 479 142.2, 2.1 HH14
141.7, 2.8 HH14
142.2, 2 0.43 WISE
FRB20211001B 143.5, -1 378.7 143.8, -1.6 HH14
S203 143.6, -1.5 3.39 ± 0.25 0.36 WISE
FRB20230709B 146.6, 4.2 514.9 146.1, 3.1 0.69 WISE
FRB20200717B 150.7, 2.9 295.2 S207 151.2, 2.5 0.46 WISE
FRB20230418E 152.3, 2.6 0.592 ± 0.59 497.5 S210 152.7, 2.9 1.67 ± 0.34 0.31 WISE
FRB20210925B 152.8, -20.9 1027 158.1, -21.4 HH14
FRB20210813B 154.4, -20.9 518.4 158.3, -20.5 HH14
158.1, -21.4 HH14
FRB20220518E 156, -17.8 377 158.3, -20.5 HH14
158.1, -21.4 HH14
FRB20200112E 156.1, -0.6 913.7 155.4, -0.3 0.53 WISE
FRB20230730A 156.3, -19.7 312.6 158.3, -20.5 HH14
158.1, -21.4 HH14
FRB20210920D 156.5, -20.5 0.618 ± 0.62 1201 158.3, -20.5 HH14
158.1, -21.4 HH14
FRB20211013D 158.4, -17.6 1091 158.3, -20.5 HH14
158.1, -21.4 HH14
FRB20220424A 158.5, -26.1 1373 158.1, -21.4 HH14
FRB20211231B 158.6, -11.9 905.7 160.5, -12.1 HH14
California Nebula; S220 160, -12.7 0.77 ± 0.31 1.58 WISE
FRB20211120B 159, -18.7 359.1 158.3, -20.5 HH14
158.1, -21.4 HH14
159.9, -18.6 0.24 ± 0.02 0.78 WISE
FRB20221012E 159.1, -15.1 0.457 ± 0.46 1118 California Nebula; S220 160, -12.7 0.77 ± 0.31 1.58 WISE
FRB20220304A 159.3, -24.2 334.3 158.1, -21.4 HH14
158.3, -20.5 HH14
FRB20210210A 159.6, -26.1 401.9 158.1, -21.4 HH14
FRB20201030C 159.7, -21.3 974.2 158.1, -21.4 HH14
158.3, -20.5 HH14
FRB20201017C 159.7, -11.3 0.659 ± 0.66 731.9 160.5, -12.1 HH14
California Nebula; S220 160, -12.7 0.77 ± 0.31 1.58 WISE
FRB20211116A 160.6, -25.4 775.1 158.1, -21.4 HH14
Table 2: Continued.
TNS Name (l, b) $\tau_{\rm sc}$ DM Name (l, b) D $\theta_{\rm IR}$ Catalog
(deg) (ms at 1 GHz) (pc cm-3) (deg) (kpc) (deg)
FRB20200323G 160.6, -9.7 0.212 ± 0.21 627.9 California Nebula; S220 160, -12.7 0.77 ± 0.31 1.58 WISE
FRB20191215C 161, -22 260.5 158.1, -21.4 HH14
158.3, -20.5 HH14
FRB20230910C 162.6, -19.1 0.634 ± 0.63 660.1 158.1, -21.4 HH14
158.3, -20.5 HH14
FRB20210507A 163.3, -8.4 525.8 165.4, -9 HH14
FRB20200131A 168.8, -21.1 1359 176.2, -20.9 HH14
FRB20230915E 169.6, -16.7 621.1 176.2, -20.9 HH14
FRB20220818B 170.7, -2 0.285 ± 0.28 1286 172, -2.2 HH14
FRB20210330D 171.3, -24.3 2280 176.2, -20.9 HH14
FRB20200626A 171.4, -3 1580 173, -1.3 HH14
172, -2.2 HH14
FRB20220405A 172.2, -17.6 462.4 176.2, -20.9 HH14
178.9, -20.1 HH14
FRB20220719F 172.4, -1 525.3 173, -1.3 HH14
172, -2.2 HH14
FRB20221025I 173.9, -21.1 453.3 178.9, -20.1 HH14
176.2, -20.9 HH14
FRB20190902B 173.9, 2.8 758.2 173.4, 2.4 HH14
173.4, 2.6 HH14
173.4, 3.2 HH14
173.6, 2.8 HH14
S235 173.5, 3.2 0.38 WISE
FRB20181119D 174.9, -0.1 566.4 173, -1.3 HH14
FRB20200929A 175.3, -16.2 821.4 178.9, -20.1 HH14
176.2, -20.9 HH14
FRB20210111D 175.4, -25.7 890.8 178.9, -20.1 HH14
176.2, -20.9 HH14
FRB20210412C 176.2, -16.6 463.9 178.9, -20.1 HH14
176.2, -20.9 HH14
FRB20230725F 176.3, -18.5 838.4 178.9, -20.1 HH14
176.2, -20.9 HH14
FRB20221218C 178.4, -16.1 372.1 176.2, -20.9 HH14
178.9, -20.1 HH14
FRB20190418A 179.4, -22.9 184.5 176.2, -20.9 HH14
178.9, -20.1 HH14
FRB20221223G 180.1, -22.1 618.8 176.2, -20.9 HH14
178.9, -20.1 HH14
FRB20190712A 180.2, -32.5 682.5 186.3, -34.3 HH14
FRB20201026E 181.1, -20.8 804.8 176.2, -20.9 HH14
178.9, -20.1 HH14
FRB20200412A 181.3, -18.3 839.5 176.2, -20.9 HH14
178.9, -20.1 HH14
FRB20230717C 181.8, -18.1 0.473 ± 0.47 231.2 178.9, -20.1 HH14
176.2, -20.9 HH14
Table 2: Continued.
TNS Name (l, b) $\tau_{\rm sc}$ DM Name (l, b) D $\theta_{\rm IR}$ Catalog
(deg) (ms at 1 GHz) (pc cm-3) (deg) (kpc) (deg)
FRB20230705E 183.2, -31 0.291 ± 0.29 596.1 186.3, -34.3 HH14
FRB20230106B 184.1, -24.3 595.7 178.9, -20.1 HH14
FRB20220710B 184.5, -34.8 190.7 186.3, -34.3 HH14
FRB20200922C 185, -29.6 498.2 186.3, -34.3 HH14
FRB20200119D 185.3, -24.4 944 178.9, -20.1 HH14
FRB20230730C 185.4, -19.4 227.2 178.9, -20.1 HH14
FRB20210327F 185.5, -29 0.293 ± 0.29 287.5 186.3, -34.3 HH14
FRB20190617B 186.2, -39.4 0.437 ± 0.44 272.6 186.3, -34.3 HH14
FRB20221021C 186.3, -19.8 721.6 178.9, -20.1 HH14
FRB20190103B 191.3, 1.1 1.008 ± 1 540.7 192.3, 0.8 0.6 WISE
FRB20210304E 192.3, -3.5 1144 192.2, -3.8 HH14
FRB20230615B 193.3, -14.7 608 194.6, -15.6 HH14
FRB20190427A 194.5, -17.1 455.9 194.6, -15.6 HH14
FRB20190409A 195.7, -17.2 1792 194.6, -15.6 HH14
FRB20230905D 196.8, -0.5 2.163 ± 2.2 443.7 S268 196.8, -2.4 1.06 WISE
FRB20191014B 196.9, -12.1 278.2 195.1, -12 HH14
S264; lambda Ori 195.3, -12.1 0.84 WISE
FRB20211204B 217.2, 1.8 425.9 216.7, 1.1 0.64 WISE

### v2

@@ VI Conclusion
Our analysis shows that FRB detections in CHIME/FRB Catalog 2 exhibit a two-dimensional dependence on Galactic coordinates. The most striking feature is the complete absence of CHIME/FRB detections in a patch of the sky that coincides with the Cygnus X region. The analyses discussed in §IIIV demonstrate that this detection gap, along with other regions of reduced FRB detections within the Galactic plane, corresponds with known H II regions. Based on the discussion presented in §V, we favor the interpretation that turbulence along these lines of sight induces large degrees of scatter-broadening that totally smear FRBs out of detectability within the CHIME band.
As shown in Section IV, measurements of $\tau_{\rm sc}$ in CHIME/FRB Catalog 2 and derived $\tau_{\rm sc}$ from non-detections can be used to gauge the accuracy of existing $n_{e}$ models of the Milky Way Galaxy. We specifically used the CHIME/FRB Catalog 2 measurements of $\tau_{\rm sc}$ to argue that the lower limit on scattering in the direction of Cygnus X is $\tau_{\rm sc,1GHz}\geq 5.59$ ms in order to produce zero detections in the CHIME band. This estimate is inconsistent with the NE2001 prediction for the maximum $\tau_{\rm sc}$; and while consistent with the YMW16 prediction, this latter model does not explicitly account for individual H II regions. Our analysis of FRB $\tau_{\rm sc}$ measurements therefore demonstrates that FRB measurements can be used to place “background" constraints on inhomogeneous $n_{e}$ models in a manner that complements the “foreground" constraints placed by radio pulsars. This prospect of using FRB $\tau_{\rm sc}$ measurements as tools to constrain $n_{e}$ models is especially important in an era where cosmological applications of FRBs are often sensitive to knowledge and/or reasonable assumptions of the $n_{e}$ distribution for both the Milky Way and FRB-host galaxies (e.g., zkz25).
@@
Future CHIME/FRB measurements of $\tau_{\rm sc}$ will increase the surface density of FRB detections at low frequencies. Such an increase will likely yield additional detection gaps associated with Galactic H II regions that have smaller angular extents. These same data sets will also allow for data-driven constraints on the Galactic electron density models themselves , through all-sky estimates of the maximum $\tau_{\rm sc}$ as estimated for the Cygnus X region in §IV.5. Although the nature of FRB classification makes similar constraints on DM more complicated, phenomenological models can be used to gauge the likely maximum values for the contribution of the Milky Way to DM (e.g., cbg+23). We therefore expect that future CHIME/FRB catalogs will provide the means to strengthen and expand upon the inferences made in our work.
@@ ACKNOWLEDGEMENTS
We are grateful to the anonymous referee for their constructive comments that have improved our work. We thank Loren Anderson, Liam Connor, Duncan Lorimer, Namir Kassim, and Jackson Taylor for thoughtful questions and discussions that informed the analyses presented in this paper.
We acknowledge that CHIME is located on the traditional, ancestral, and unceded territory of the Syilx/Okanagan people. We are grateful to the staff of the Dominion Radio Astrophysical Observatory, which is operated by the National Research Council of Canada. CHIME operations are funded by a grant from the NSERC Alliance Program and by support from McGill University, University of British Columbia, and University of Toronto. CHIME was funded by a grant from the Canada Foundation for Innovation (CFI) 2012 Leading Edge Fund (Project 31170) and by contributions from the provinces of British Columbia, Québec and Ontario. The CHIME/FRB Project was funded by a grant from the CFI 2015 Innovation Fund (Project 33213) and by contributions from the provinces of British Columbia and Québec, and by the Dunlap Institute for Astronomy and Astrophysics at the University of Toronto. Additional support was provided by the Canadian Institute for Advanced Research (CIFAR), the Trottier Space Institute at McGill University, and the University of British Columbia (UBC). This research has made use of the VizieR catalogue access tool, CDS, Strasbourg, France (DOI : 10.26093/cds/vizier). The original description of the VizieR service was published in 2000, A&AS 143, 23.
@@ Appendix A Literature Scattering Values near Cygnus X
Table 1: Table with literature values of scattering and angular broadening towards Cygnus X, from background pulsars, FRBs, quasars, and Cygnus X-3. References: [1] jing+25, [2] CHAMPSS+25, [3] zhz_2022, [4] fey+89, [5] fey+91, [6] CygnusX-molnar
Source $\ell$ (deg) $b$ (deg) $\theta$ (mas) $\tau_{\mathrm{1\,GHz}}$ (ms) $\tau_{\mathrm{600\,MHz}}$ (ms) DM ( $\text{ pc cm}^{-3}$) Reference
PSR J2021+4024g 78.18 2.11 8.21e+01 6.33e+02 678.8 [1]
PSR J2030+3818g 77.45 -0.51 7.80 6.02e+01 596.8 [1]
PSR J2030+3944g 78.60 0.30 1.14e+02 8.80e+02 934.8 [1]
PSR J2046+4253g 83.02 -0.26 1.27e+02 9.80e+02 622.2 [1]
PSR J2052+4421g 84.84 -0.17 9.50e+01 7.33e+02 543.8 [1]
PSR J2108+5001 91.20 1.47 1.91 14.7 482.3 [2]
FRB 20210705 76.03 2.71 1.19e+01 9.20e+01 2011.6 [3]
@@ B.1 Catalog
Table 2: A full list of FRB sightlines intersecting H II regions from the WISE (wise) catalog.
TNS Name (l, b) $\tau_{\rm sc}$ DM Name (l, b) D $\theta_{\rm IR}$ $\eta$
(deg) (ms at 1 GHz) (pc cm-3) (deg) (kpc) (deg) ($\theta_{\rm IR}$)
FRB20200306C 15.6, 28 337.2 6.4, 22.9 5.19 1.89
FRB20230511B 15.7, 23.1 191.6 6.4, 22.9 5.19 1.65
FRB20181119B 67.8, 1 2.307 ± 0.043 608.5 68.2, 1.1 0.22 1.8
FRB20230426F 99.4, 1.4 0.286 ± 0.017 1360.1 S131; IC1396 99.5, 3.8 0.91 ± 0.57 1.26 1.89
FRB20210810B 101.4, 4.4 1.101 ± 0.153 639.6 S131; IC1396 99.5, 3.8 0.91 ± 0.57 1.26 1.58
FRB20210103A 106.7, -0.2 1.881 ± 0.185 746.2 S142 107, -0.8 3.52 ± 0.47 0.37 1.83
FRB20230102D 107.2, 6 983.0 S141 107.9, 5.6 0.55 1.37
S150 108.5, 6.4 1.34 ± 0.47 0.82 1.63
FRB20230316H 117.3, 3.9 496.1 S171 118.2, 5.4 1.39 1.32
FRB20220411A 117.5, 8 0.323 ± 0.059 507.4 S171 118.2, 5.4 1.39 1.95
FRB20210309F 118.5, 6.5 547 S171 118.2, 5.4 1.39 0.82
FRB20200828D 118.9, 7.7 0.96 ± 0.052 1516.6 S171 118.2, 5.4 1.39 1.74
FRB20190912D 119.3, 3.9 786.2 S171 118.3, 4.9 0.84 ± 0.31 0.73 1.32
S171 118.2, 5.4 1.39 1.35
FRB20220218A 120.1, 5.4 594.5 S171 118.2, 5.4 1.39 1.85
FRB20220330A 124.9, 0.9 0.367 ± 0.023 664.9 125, 0.9 0.12 0.95
FRB20220814A 127.5, 1 0.919 ± 0.104 1430.7 127.6, 1.1 0.2 0.59
FRB20190302A 133.2, 0.2 11.28 ± 0.604 1033.8 S190; W3 133.5, 0.8 0.46 1.31
FRB20200416A 134.8, 1.2 0.702 ± 0.035 538.6 W4; S190 135.8, 0.9 0.69 0.6
W4; S190 135, 0.7 3.31 ± 0.35 0.87 1.46
FRB20200708A 138.2, -1.1 7.998 ± 0.667 842.5 139.7, -1.2 1.52 1.99
FRB20191209B 138.9, 2.9 910.7 139.6, 2.5 3.81 ± 0.34 0.38 1.05
FRB20221129E 140.8, 2.7 1.709 ± 0.372 2293.1 140.8, 3.1 0.21 1.9
FRB20210207E 142.6, 2 479 142.2, 2 0.43 1.08
FRB20211001B 143.5, -1 378.7 S203 143.6, -1.5 3.39 ± 0.25 0.36 1.57
FRB20230709B 146.6, 4.2 514.9 146.1, 3.1 0.69 1.75
FRB20200717B 150.7, 2.9 295.2 S207 151.2, 2.5 0.46 1.46
FRB20230418E 152.3, 2.6 0.592 ± 0.051 497.5 S210 152.7, 2.9 1.67 ± 0.34 0.31 1.65
FRB20200112E 156.1, -0.6 913.7 155.4, -0.3 0.53 1.49
FRB20211231B 158.6, -11.9 905.7 California Nebula; S220 160, -12.7 0.77 ± 0.31 1.58 1.12
FRB20211120B 159,-18.7 359.1 159.9, -18.6 0.24 ± 0.02 0.78 0.96
FRB20221012E 159.1, -15.1 0.457 ± 0.08 1118.3 California Nebula; S220 160, -12.7 0.77 ± 0.31 1.58 1.57
FRB20201017C 159.7,-11.3 0.659 ± 0.048 731.9 California Nebula; S220 160, -12.7 0.77 ± 0.31 1.58 0.91
FRB20200323G 160.6, -9.7 0.212 ± 0.05 627.9 California Nebula; S220 160, -12.7 0.77 ± 0.31 1.58 0.95
FRB20190902B 173.9, 2.8 758.2 S235 173.5, 3.2 0.38 1.54
FRB20190103B 191.3, 1.1 1.008 ± 0.121 540.7 192.3, 0.8 0.6 1.78
FRB20230905D 196.8, -0.5 2.163 ± 0.006 443.7 S268 196.8, -2.4 1.06 1.84
FRB20191014B 196.9,-12.1 278.2 S264; lambda Ori 195.3, -12.1 0.84 1.73
FRB20211204B 217.2, 1.8 425.9 216.7, 1.1 0.64 1.28